OpenStudy (anonymous):

if h(x)=2g(x) - (1)/(g(x)) =2 and g'(0)=-1, find the exact value of h'(0). please please please help and show all steps!

OpenStudy (jamesj):

what exactly is h? h = 2g - (1/g); or h = (2g - 1)/g or something else?

OpenStudy (anonymous):

h is 2g - (1/g)

OpenStudy (jamesj):

Ok. So using the chain rule as I showed you earlier, what is h'? dh/dx = dh/dg . dg/dx = ... what?

OpenStudy (anonymous):

okay is the answer 4?

OpenStudy (jamesj):

what is g(0) equal to? is it the = 2 in your question?

OpenStudy (anonymous):

yeah sorry g(0)=2

OpenStudy (jamesj):

then the answer is not 4. What's your expression for h'(x) ?

OpenStudy (anonymous):

h'(x)=2 + (g'(x)/(g(x)^2))

OpenStudy (jamesj):

No. dh/dx = dh/dg . dg/dx = (2 + 1/g^2) g' i.e., h'(x) = g'(x) [ 2 + 1/g(x)^2 ]

OpenStudy (jamesj):

yes?

OpenStudy (anonymous):

okay so is it -9/4?

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

okay thank you!